# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ 4 Asymptotic properties of ABC with maximum mean discrepancy
Although the theoretical results in Section 24 provide an improved understanding of the limiting properties of discrepancy–based abc posteriors, the i.i.d. assumption in (I) rules out important settings which typically require abc implementations. A remarkable case is that of time–dependent observations (e.g., Fearnhead & Prangle, 2012; Bernton et al., 2019; Nguyen et al., 2020; Drovandi & Frazier, 2022). Section 6.1 clarifies that the theory derived under i.i.d.  assumptions in Section 24 can be naturally extended to these non–i.i.d. settings leveraging results for the Rademacher complexity of $\beta$–mixing stochastic processes (Mohri & Rostamizadeh, 2008).
6.1 Convergence and concentration beyond i.i.d. settings
@@
In the following, the above product structure is not imposed. Instead, we only assume that the marginal of $\mu^{*(n)}$ is constant, and is denoted with $\mu^{*}$. Such an assumption is met whenever $y_{1:n}$ is extracted from a stationary process $(y_{t})_{t\in\mathbb{Z}}$, thus embracing a broader variety of applications of direct interest. Under these settings, a statistical model is defined as a collection of distributions in $\mathcal{P}(\mathcal{Y}^{n})$, i.e., $\{\mu_{\theta}^{(n)}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, with a constant marginal distribution denoted by $\mu_{\theta}$. Notice that these assumptions of constant marginals $\mu^{*}$ and $\mu_{\theta}$ are made also in available concentration theory under non–i.i.d. settings (Bernton et al., 2019; Nguyen et al., 2020) when requiring convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and suitable concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$. As such, the settings we consider are not more restrictive than those addressed in discrepancy–specific theory. In fact, both Bernton et al. (2019) and Nguyen et al. (2020) explicitly refer to stationary processes when discussing the validity of the assumptions on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ in non–i.i.d. contexts.
@@
$\pi_{n}^{(\varepsilon_{n})}(\theta)\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(z_{1:n},y_{1:n})\leq\varepsilon_{n}\}\ \mu_{\theta}^{(n)}(dz_{1:n}).$
Notice that the above definition is the same as the one provided in Section 2, with the only difference that $\mu_{\theta}^{n}=\prod_{i=1}^{n}\mu_{\theta}$ is replaced by the joint distribution $\mu^{(n)}_{\theta}$, since in this case the data are no more assumed to be independent.
In order to extend the convergence result in Corollary 3.1 together with the concentration statement in Theorem 3.2 to the above framework, we require an analog of equation (2.2) in Lemma 2.1 for time–dependent data. This generalization can be derived leveraging results in Mohri & Rostamizadeh (2008) under the notion of $\beta$–mixing coefficients.
Definition 6.1 ($\beta$–mixing; e.g., Mohri & Rostamizadeh (2008)).
Consider a stationary sequence $(x_{t})_{t\in\mathbb{Z}}$ of random variables, and let $\sigma^{j^{\prime}}_{j}$ be the $\sigma$–algebra generated by the random variables $x_{k}$, $j\leq k\leq j^{\prime}$, for any $j,j^{\prime}\in\mathbb{Z}\cup\{-\infty,+\infty\}$. Then, for any positive integer $k$, the $\beta$–mixing coefficient of $(x_{t})_{t\in\mathbb{Z}}$ is defined as
@@
If $\beta(k)\to 0$ as $k\to\infty$, then the stochastic process $(x_{t})_{t\in\mathbb{Z}}$ is said to be $\beta$–mixing.
Intuitively, $\beta(k)$ measures the dependence between the past (before $t$) and the future (after $t+k$) of the process. When such a dependence is weak, we expect that $\beta(k)$ will decay to 00 fast when $k\rightarrow\infty$. In the most extreme case, when the $x_{t}$’s are i.i.d., we have $\beta(k)=0$ for all $k>0$. More generally, as clarified in Definition 6.1, a process having $\beta(k)\rightarrow 0$ when $k\rightarrow\infty$ is named $\beta$–mixing. We refer the reader to Doukhan (1994) for an in–depth study of the main properties of $\beta$–mixing processes along with a more comprehensive discussion of relevant examples. The most remarkable ones will be also presented in the following.
Leveraging the above notion of $\beta$–mixing coefficient, Lemma 6.1 extends Lemma 2.1 to the dependent setting. The proof of Lemma 6.1 can be found in the Supplementary Material and combines Proposition 2 and Lemma 2 in Mohri & Rostamizadeh (2008). For readability, let us also introduce the notation $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Note that $s_{n}\sim\sqrt{n/2}$ as $n\to\infty$, and thus $s_{n}\to\infty$.
Lemma 6.1.
Define $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Moreover, consider the stationary stochastic process $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\beta(k),k\in\mathbb{N}$, its $\beta$–mixing coefficients. Let $\mu^{(n)}$ be the joint distribution of a sample $x_{1:n}$ extracted from $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\mu=\mu^{(1)}$ its constant marginal. Then, for any $b$–uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and scalar $\delta\geq 0$,
$\qquad\ \ \ \mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\right]\geq 1-2{\exp}\left[-\frac{s_{n}\delta^{2}}{2b^{2}}\right]-2s_{n}\beta({\lfloor}\sqrt{n}{\rfloor}),$ (6.1)
where $\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})$ corresponds to the Rademacher complexity in Definition 2.2 for an i.i.d. sample of size $s_{n}$ from $\mu$.
Equation (6.1) extends (2.2) beyond the i.i.d. setting. This extension crucially provides a bound that still depends on the Rademacher complexity introduced in Definition 2.2 for an i.i.d. sample — in this case from the common marginal $\mu$ of the stochastic process $(x_{t})_{t\in\mathbb{Z}}$. As such, Assumption (IV) requires no modifications and, hence, no additional validity checks relative to those discussed in Section 3.3. This result suggests that the Rademacher complexity framework might also be leveraged to derive improved convergence and concentration results for discrepancy–based abc posteriors in more general situations which do not necessarily meet Assumption (I). To prove these results we leverage again Assumptions (II), (III) and (IV), and replace (I) with condition (VII) below.
@@
    The data $y_{1:n}$ are from a $\beta$–mixing stochastic process $(y_{t})_{t\in\mathbb{Z}}$ with mixing coefficients $\beta(k)\leq C_{\beta}e^{-\gamma k^{\xi}}$ for some $C_{\beta},\gamma,\xi>0$, common marginal $\mu$, and generic joint $\mu^{*(n^{\prime})}$ for a sample $y_{1:n^{\prime}}$ from $(y_{t})_{t\in\mathbb{Z}}$ for any $n^{\prime}\in\mathbb{N}$. The same $\beta$–mixing conditions hold also for the stochastic process $(z_{t})_{t\in\mathbb{Z}}$ associated with the synthetic data $z_{1:n}$ generated from the assumed model. In this case, the joint distribution for a generic sample $z_{1:n^{\prime}}$ is $\mu^{(n^{\prime})}_{\theta}$, $\theta\in\Theta$, whereas the common marginal is denoted by $\mu_{\theta}$.
@@
Assumption (VII) is clearly more general than (I). As discussed previously, it embraces several stochastic processes of substantial interest in practical applications, including those in Examples 6.16.2 below; see Doukhan (1994) for additional examples and discussion.
Example 6.1 (Doeblin–recurrent Markov chains).
@@
Let $(x_{t})_{t\in\mathbb{Z}}$ be a Markov chain on $\mathcal{Y}\subset\mathbb{R}^{d}$ with transition kernel $P(\cdot,\cdot)$. Such a Markov chain is said to be Doeblin–recurrent if there exists a probability measure $q$, a constant $10<c<$ 1 and an integer $r>0$ such that, for any measurable set $A$ and any $x\in\mathbb{R}^{d}$, $P^{r}(x,A)>cq(A)$. When this is the case, $(x_{t})_{t\in\mathbb{Z}}$ is $\beta$–mixing with $\beta(k)\leq 2(1-c)^{k/r}$; see e.g., Theorem 1 in page 88 of Doukhan (1994).
Example 6.2 (Hidden Markov chains).
@@
Assume $(x_{t})_{t\in\mathbb{Z}}$ is a $\beta$–mixing stochastic process with coefficients $\beta_{x}(k),k\in\mathbb{N}$. If $\tilde{x}_{t}=F(x_{t},\varepsilon_{t})$ with $\varepsilon_{t}$ i.i.d., then the $\beta$–mixing coefficients of $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ satisfy $\beta_{\tilde{x}}(k)=\beta_{x}(k)$. Therefore, $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is also $\beta$–mixing and inherits the bounds on $\beta_{x}(k)$. These processes are often used in practice with $(x_{t})_{t\in\mathbb{Z}}$ being a Markov chain. In this case $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is called a Hidden Markov chain.
@@
Section 2.4.2 of Doukhan (1994) also provides conditions on $F$ and on the i.i.d. sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ ensuring that a stationary process $(x_{t})_{t\in\mathbb{Z}}$ satisfying $x_{t}=F(x_{t-1},\dots,x_{t-k},\varepsilon_{t})$ exists and is $\beta$–mixing. Lemma 6.2 specializes such a result in the context of Gaussian ar(1) processes, which will be considered in the empirical study in Section 7.
Lemma 6.2 (Gaussian ar(1) process).
@@
Consider a generic sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ of i.i.d random variables from a $\mbox{N}(0,\sigma^{2})$. Moreover, let $-1<\theta<1$ and $\psi\in\mathbb{R}$. Then the stationary solution to $x_{t}=\psi+\theta x_{t-1}+\varepsilon_{t}$ is $\beta$–mixing and has coefficients $\beta(k)\leq\smash{|\theta|^{k}/(2\sqrt{1-\theta^{2}})}=(2\sqrt{1-\theta^{2}})^{-1}\exp(-k\log(1/|\theta|))$, $k\in\mathbb{N}$, thus meeting (VII).
@@
Note that in the empirical study in Section 7 the focus will be on Bayesian inference for the parameter $\theta$. Clearly, in this case it is not sufficient to focus on the marginal distribution of each $x_{t}$. Rather, one should leverage the bivariate distribution for the pairs $\tilde{x}_{t}:=(x_{t},x_{t+1})$; see also Bernton et al. (2019) where such a strategy is named delay reconstruction. This procedure simply changes the focus to the bivariate stochastic process $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, but does not alter the mixing properties. In particular, if $\beta_{x}(k)$ and $\beta_{\tilde{x}}(k)$ are the mixing coefficients of $(x_{t})_{t\in\mathbb{Z}}$ and $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, respectively, then from Definition 6.1 we have $\beta_{\tilde{x}}(k)=\beta_{x}(k-1)$, for $k\geq 1$.
@@
Leveraging Lemma 6.1 along with the newly–introduced assumption, Proposition 6.1 states convergence of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\to\infty$.
Proposition 6.1.
@@
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (6.2)
@@
According to Proposition 6.1, replacing Assumption (I) with (VII), does not alter the uniform convergence properties of the abc posterior originally stated in Corollary 3.1 under the i.i.d. assumption. This allows to inherit the discussion after Corollary 3.1 also beyond i.i.d. settings, while suggesting that similar extensions would be possible in the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$. These extensions are stated in Theorem 6.1, which provides an important generalization of Theorem 3.2 beyond the i.i.d. case.
Theorem 6.1.
@@
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\left[\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\right]^{1/2}\Bigr{\}}\Bigr{)}\leq\frac{4\cdot 3^{L}}{c_{\pi}n},$

### v5

@@ Appendix B Concentration in the space of parameters
Although the theoretical results in Sections 24 provide an improved understanding of the limiting properties of discrepancy-based abc posteriors, the i.i.d. assumption in (I) rules out important settings which often require abc. A remarkable case is that of time-dependent observations (e.g., Fearnhead and Prangle, 2012; Bernton et al., 2019; Nguyen et al., 2020; Drovandi and Frazier, 2022).
Section C.1 clarifies that the theory derived under i.i.d.  assumptions in Section 24 can be naturally extended to these non-i.i.d. settings leveraging results for Rademacher complexity in $\beta$-mixing stochastic processes (Mohri and Rostamizadeh, 2008). Examples C.3C.4 below show that such a class embraces several processes of direct practical interest. Extensions beyond this class, albeit relevant, are challenging even when the focus is on proving simpler, non-uniform, concentration results for a single discrepancy. Hence, these extensions are left for future research, that could be facilitated by the derivation of Rademacher complexity bounds for general processes beyond the $\beta$-mixing ones studied in Mohri and Rostamizadeh (2008).
C.1 Convergence and concentration beyond i.i.d. settings
@@
In the following, the above product structure is not imposed. Instead, we only assume that the marginal of $\mu^{*(n)}$ is constant, and denoted with $\mu^{*}$. Such an assumption is met whenever $y_{1:n}$ is extracted from a stationary stochastic process $(y_{t})_{t\in\mathbb{Z}}$, thus embracing a broader variety of applications of direct interest. Under these settings, a statistical model is defined as a collection of distributions in $\mathcal{P}(\mathcal{Y}^{n})$, i.e., $\{\mu_{\theta}^{(n)}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, with a constant marginal denoted by $\mu_{\theta}$. Notice that these assumptions of constant marginals $\mu^{*}$ and $\mu_{\theta}$ are made also in the available concentration theory under non-i.i.d. settings (see e.g., Bernton et al., 2019; Nguyen et al., 2020) when requiring convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and suitable concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$. As a result, the settings we consider are not more restrictive than those addressed in discrepancy-specific theory. In fact, both Bernton et al. (2019) and Nguyen et al. (2020) explicitly refer to stationary processes when discussing the validity of the assumptions on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ in non-i.i.d. contexts.
@@
$\pi_{n}^{(\varepsilon_{n})}(\theta)\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon_{n}\}\ \mu_{\theta}^{(n)}(dz_{1:n}).$
This definition is the same as the one in Section 2, with the only difference that $\mu_{\theta}^{n}=\prod_{i=1}^{n}\mu_{\theta}$ is replaced by the joint $\mu^{(n)}_{\theta}$, since now the data are no more assumed to be independent.
In order to extend the convergence result in Corollary 3.2 together with the concentration statement in Theorem 3.3 to the above framework, we require an analog of Equation (2) in Lemma 2.6 for time-dependent data. This generalization can be derived leveraging results in Mohri and Rostamizadeh (2008) under the notion of $\beta$-mixing coefficients.
Definition C.1 ($\beta$-mixing).
Consider the stationary sequence $(x_{t})_{t\in\mathbb{Z}}$ of random variables, and let $\sigma^{j^{\prime}}_{j}$ be the $\sigma$–algebra generated by the random variables $x_{k}$, $j\leq k\leq j^{\prime}$, for any $j,j^{\prime}\in\mathbb{Z}\cup\{-\infty,+\infty\}$. Then, for any integer $k>0$, the $\beta$-mixing coefficient of $(x_{t})_{t\in\mathbb{Z}}$ is defined as
@@
If $\beta(k)\to 0$ as $k\to\infty$, then the stochastic process $(x_{t})_{t\in\mathbb{Z}}$ is said to be $\beta$-mixing.
Intuitively, $\beta(k)$ measures the dependence between the past (before $t$) and the future (after $t+k$) of the process. When such a dependence is weak, we expect that $\beta(k)$ will decay to 00 fast when $k\rightarrow\infty$. In the most extreme case, when the $x_{t}$’s are i.i.d., we have $\beta(k)=0$ for all $k>0$. More generally, as clarified in Definition C.1, a process having $\beta(k)\rightarrow 0$ when $k\rightarrow\infty$ is named $\beta$-mixing. We refer the reader to Doukhan (1994) for an in-depth study of the main properties of $\beta$-mixing processes along with a more comprehensive discussion of relevant examples. The most remarkable ones will be also presented in the following.
Leveraging the notion of $\beta$-mixing coefficient, Lemma C.2 extends Lemma 2.6 to the dependent setting. The proof can be found in Appendix D and combines Proposition 2 and Lemma 2 in Mohri and Rostamizadeh (2008). For readability, let us also introduce the notation $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Note that $s_{n}\sim\sqrt{n}/2$ as $n\to\infty$, and thus $s_{n}\to\infty$.
Lemma C.2.
Define $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Moreover, consider the stationary stochastic process $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\beta(k),k\in\mathbb{N}$, its $\beta$-mixing coefficients. Let $\mu^{(n)}$ be the joint distribution of a sample $x_{1:n}$ extracted from $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\mu=\mu^{(1)}$ its constant marginal. Then, for any $b$-uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and scalar $\delta\geq 0$,
$\qquad\ \ \ \mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\right]\geq 1-2{\cdot}{\exp}\left[-\frac{s_{n}\delta^{2}}{2b^{2}}\right]-2s_{n}\beta({\lfloor}\sqrt{n}{\rfloor}),$ (C.1)
with $\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})$ the Rademacher complexity in Definition 2.5 for an i.i.d. sample of size $s_{n}$ from $\mu$.
Equation (C.1) extends (2) beyond i.i.d. settings. This extension provides a bound that still depends on the Rademacher complexity in Definition 2.5 for an i.i.d. sample — in this case from the common marginal $\mu$ of the process $(x_{t})_{t\in\mathbb{Z}}$. As such, Assumption (IV) requires no modifications, and no additional validity checks relative to those discussed in Section 3.3. This suggests that the Rademacher complexity framework might also be leveraged to derive improved convergence and concentration results for discrepancy-based abc posteriors in more general situations which do not necessarily meet Assumption (I). To prove these results we leverage again Assumptions (II), (III) and (IV), and replace (I) with condition (VII).
@@
    The data $y_{1:n}$ are from a $\beta$-mixing stochastic process $(y_{t})_{t\in\mathbb{Z}}$ with mixing coefficients $\beta(k)\leq C_{\beta}e^{-\gamma k^{\xi}}$ for some $C_{\beta},\gamma,\xi>0$, common marginal $\mu$, and generic joint $\mu^{*(n^{\prime})}$ for a sample $y_{1:n^{\prime}}$ from $(y_{t})_{t\in\mathbb{Z}}$ for any $n^{\prime}\in\mathbb{N}$. The same $\beta$-mixing conditions hold also for the process $(z_{t})_{t\in\mathbb{Z}}$ associated with the synthetic data $z_{1:n}$ from the assumed model. In this case, the joint distribution for a generic sample $z_{1:n^{\prime}}$ is $\mu^{(n^{\prime})}_{\theta}$$\theta\in\Theta$, and the common marginal is denoted by $\mu_{\theta}$. For simplicity and without loss of generality, we also assume that the constants $C_{\beta}$, $\gamma$, and $\xi$ are the same for $(y_{t})_{t\in\mathbb{Z}}$ and $(z_{t})_{t\in\mathbb{Z}}$.
@@
Assumption (VII) is clearly more general than (I). As discussed previously, it embraces several stochastic processes of substantial interest in practical applications, including those in Examples C.3C.4 below; see Doukhan (1994) for additional examples and discussion.
Example C.3 (Doeblin-recurrent Markov chains).
@@
Let $(x_{t})_{t\in\mathbb{Z}}$ be a Markov chain on $\mathcal{Y}\subset\mathbb{R}^{d}$ with transition kernel $P(\cdot,\cdot)$. Such a Markov chain is said to be Doeblin-recurrent if there exists a probability measure $q$, a constant $0<c\leq 1$ and an integer $r>0$ such that, for any measurable set $A$ and any $x\in\mathbb{R}^{d}$, $P^{r}(x,A)\geq cq(A)$. When this is the case, $(x_{t})_{t\in\mathbb{Z}}$ is $\beta$-mixing with $\beta(k)\leq 2(1-c)^{k/r}$; see e.g., Theorem 1 in page 88 of Doukhan (1994).
Example C.4 (Hidden Markov chains).
@@
Assume $(x_{t})_{t\in\mathbb{Z}}$ is a $\beta$-mixing stochastic process with coefficients $\beta_{x}(k),k\in\mathbb{N}$. If $\tilde{x}_{t}=F(x_{t},\varepsilon_{t})$ with $\varepsilon_{t}$ i.i.d., then the $\beta$-mixing coefficients of $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ satisfy $\beta_{\tilde{x}}(k)=\beta_{x}(k)$. Therefore, $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is also $\beta$-mixing and inherits the bounds on $\beta_{x}(k)$. These processes are often used in practice with $(x_{t})_{t\in\mathbb{Z}}$ being a Markov chain. In this case $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is called a Hidden Markov chain.
@@
Section 2.4.2 of Doukhan (1994) also provides conditions on $F$ and on the i.i.d. sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ ensuring that a stationary process $(x_{t})_{t\in\mathbb{Z}}$ satisfying $x_{t}=F(x_{t-1},\dots,x_{t-k},\varepsilon_{t})$ exists and is $\beta$-mixing. Lemma C.5 specializes such a result in the context of Gaussian ar(1) processes, which will be considered in the empirical study in Section C.2.
Lemma C.5 (Gaussian ar(1) process).
@@
Consider a generic sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ of i.i.d. random variables from a $\mbox{N}(0,\sigma^{2})$. Moreover, let $-1<\theta<1$ and $\psi\in\mathbb{R}$. Then the stationary solution to $x_{t}=\psi+\theta x_{t-1}+\varepsilon_{t}$ is $\beta$-mixing and has coefficients $\beta(k)\leq\smash{|\theta|^{k}/(2\sqrt{1-\theta^{2}})}=(2\sqrt{1-\theta^{2}})^{-1}\exp(-k\log(1/|\theta|))$, $k\in\mathbb{N}$, thus meeting (VII).
@@
Notice that in the empirical study in Section C.2 the focus will be on inference for the ar parameter $\theta$. Clearly, in this case it is not sufficient to focus on the marginal distribution of each $x_{t}$. Rather, one should leverage the bivariate distribution for the pairs $\tilde{x}_{t}:=(x_{t},x_{t+1})$; see also Bernton et al. (2019) where such a strategy is named delay reconstruction. This procedure simply changes the focus to the bivariate stochastic process $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, but does not alter the mixing properties. In particular, if $\beta_{x}(k)$ and $\beta_{\tilde{x}}(k)$ are the mixing coefficients of $(x_{t})_{t\in\mathbb{Z}}$ and $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, respectively, then from Definition C.1 we have $\beta_{\tilde{x}}(k)=\beta_{x}(k-1)$, for $k\geq 1$. Notice that identifiability is a key to ensure concentration in the space of parameters of interest as in Corollary B.1. This motivates further research, beyond the scope of this article, to derive delay reconstruction strategies ensuring identifiability in more complex processes.
@@
Leveraging Lemma C.2 along with the newly-introduced Assumption (VII), Proposition C.6 states convergence of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\to\infty$.
Proposition C.6.
@@
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (C.2)
@@
According to Proposition C.6, replacing Assumption (I) with (VII), does not alter the uniform convergence properties of the abc posterior originally stated in Corollary 3.2 under the i.i.d. assumption. This allows to inherit the discussion after Corollary 3.2 also beyond i.i.d. settings, while suggesting that similar extensions would be possible in the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$. These extensions are stated in Theorem C.7, which provides an important generalization of Theorem 3.3 beyond the i.i.d. case.
Theorem C.7.
@@
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}\Bigr{\}}\Bigr{)}\leq\frac{4\cdot 3^{L}}{c_{\pi}n},$